Β Eliminate π‘Ž, 𝑏 and 𝑐 from 𝑧 = π‘Ž(π‘₯ + 𝑦) + 𝑏(π‘₯ + 𝑦) + π‘Žπ‘π‘‘ + 𝑐.


1
Expert's answer
2021-07-26T13:15:07-0400

Eliminate π‘Ž, 𝑏 and 𝑐 from "\ud835\udc67 = \ud835\udc4e(\ud835\udc65 + \ud835\udc66) + \ud835\udc4f(\ud835\udc65 - \ud835\udc66) + \ud835\udc4e\ud835\udc4f\ud835\udc61 + \ud835\udc50."

Differentiate the given equation partially with respect to "x, y," and "t"


"\dfrac{\partial z}{\partial x}=a+b, \dfrac{\partial z}{\partial y}=a-b, \dfrac{\partial z}{\partial t}=ab"

"(a+b)^2-(a-b)^2=a^2+2ab+b^2-a^2+2ab-b^2=4ab"

Then


"(\dfrac{\partial z}{\partial x})^2-(\dfrac{\partial z}{\partial y})^2=4\dfrac{\partial z}{\partial t}"

This is a partial differential equation of order one and degree two.